OpenStudy (anonymous):

find the equation of a circle radius is 2 and through (1,2) and (3,4)

OpenStudy (lgbasallote):

hint \[\sqrt{(3-1)^2 + (4-2)^2} = 2r\]

OpenStudy (anonymous):

Equation of Circle is given by: \[\large (x-h)^2 + (y-k)^2 = r^2\]

OpenStudy (anonymous):

i started by equating 2 to the distance formula of (1,2) and (3,4) ---2 equations but then i'm now stuck because i got h+k=5

OpenStudy (anonymous):

let the center by (Cx, Cy)

OpenStudy (anonymous):